An inclined plane reduces the effort force by
1
Expert's answer
2020-04-02T10:56:22-0400

An inclined plane reduces the effort force by increasing the distance through which the force is applied.

And the work is defined as scalar product of force and displacement:


"W=Fs\cos{\alpha}\leq Fs"

So the actual effort decreases and is less than the unity always.

Therefore, an inclined plane reduces the effort as well as work done.


Answer: increasing the distance.